# thanks: ISTI Fellow

### v1

@@ III.2 Allowing no uncertainty in ejecta velocity ($\beta_{\phi}=0$)
However, without an additional parameter to encapsulate the ejecta systematics, we find several notable differences, largely dictated by the $v_{\rm{ej}}$ contours in Figure 1. Specifically, we find that for all three ejecta fits, the GW+EM analyses all favor large radii in order to replicate the low dynamical ejecta velocities necessitated by the $v_{\rm{ej}}=v_{d}$ posteriors in Figure 3. Based on Figure 1, a tight constraint on $v_{\rm ej}$ requires either asymmetric NS at more conventional radius ($\delta\simeq 0.4$ and $R_{1.4}\simeq 15{\rm km}$) or symmetric NS with extreme radius. Indeed, the union of all three $\delta-R_{1.4}$ posteriors in Figure 6 closely mirrors the underlying $v_{\rm ej}$ contours. As a result, this constraint forces the posterior away from the peak GW likelihood, which favors lower NS radii and more symmetric binaries. In all three cases where we omit systematic error in velocity ($\beta_{\phi}=0$), joint GW+EM inference mildly favors asymmetric binaries $\delta\simeq 0.1-0.2$ with large NS radius $R_{1.4}>15{\rm km}$.
III.3 Allowing no uncertainty in ejecta velocity and disk mass ejection fraction ($\beta_{\phi}$ = 0 and $f_{\rm{disk}}=0.30$)
We also consider the case of a fixed disk mass ejection fraction, in which $f_{\rm{disk}}$ is not a free parameter, but restricted to the fixed value $f_{\rm{disk}}$ $=30\%$. The purpose of this analysis is to evaluate the behavior of the ejecta fits in predicting light curves under the assumption of a generally applied ejected disk fraction. The top half of Figure 7 displays the same ejecta parameters as in Figure 3. While the dynamical mass $m_{\rm{d}}$ and viewing angle $\theta$ are somewhat consistent with prior results, the wind mass $m_{\rm{w}}$ and dynamical velocity $v_{\rm{d}}$ posteriors yield substantially different results. In the bottom half of Figure 7, we plot the light curves corresponding to the Nedora21 GW+EM posteriors shown in the top part of Figure 7. While the light curves corresponding to the median posterior values (solid lines) are clearly worse fits than those in Figure 4, the more striking feature lies in the exceptionally large uncertainties across all broadband filters. More specifically, the uncertainty bounds are much broader than before, resulting from the broad dynamic range of the $m_{\rm{w}}$ posteriors. The difficulty in constraining $m_{\rm{w}}$ comes from $f_{\rm{disk}}$ being directly tied to the amount of ejected wind mass; as such, what the $m_{w}$ posteriors in Figure 7 are really showing us is the Monte Carlo integrator trying to reconcile the values of $m_{w}$ that actually fit the data with the constricting value of $f_{\rm{disk}}$ $=30\%$ that fixes the wind ejecta mass.
@@ IV Discussion
Unsurprisingly, inference analyses using a joint EM and GW dataset are more constraining than using EM-only data. This behavior is most evident for the $\mathcal{M}_{c}$ and $\delta$ posteriors in Figure 5, which highlight the impact of including the gravitational likelihood $\mathcal{L}_{\rm{GW}}$. The $\mathcal{M}_{c}$ posteriors for each fit clearly converge on a preferred value when supplemented with GW information; likewise, the $\delta$ posteriors become narrower on both global (dynamic range of posterior values) and local (per-ejecta-fit posterior values) scales. With regard to the NS radius $R_{\rm{1.4}}$ posteriors, the difference between EM and GW+EM analyses is less striking. More prevalent, however, are the different regions of the NS radius parameter space covered by the posteriors corresponding to each ejecta fit. Specifically, the GW+EM predictions for $R_{\rm{1.4}}$ are roughly between 8–15 km, 10–20 km, and 10–40 km for the Nedora21  KruFo20  and DiCo20 forward models, respectively. These results strongly imply that, provided a constraining kilonova model, multimessenger inference about neutron star radius and, consequently, the dense matter equation of state is entirely dependent on the chosen forward model.
@@
Keeping the relative impact of GW and EM information in mind, the joint GW/EM inferences shown in Figure 5 make sense. For almost all parameters, the addition of GW information only slightly perturbs the answers derived from EM information alone. This behavior is most evident in the bottom panel of Figure 5, where narrower priors on $f_{\rm{disk}}$ directly affect the ejected wind mass and thus the EM likelihood, shifting the $\delta$ and $R_{\rm{1.4}}$ posteriors to better agreement across ejecta fits.
@@
As explained in Section II.1, we introduce the $\alpha_{\rm{dyn}}$ , $f_{\rm{disk}}$ , and $\beta_{\phi}$ free parameters to quantify the systematic uncertainty of the ejecta fit predictions, with $\sigma_{\rm{sys}}$ encompassing all other sources of systematic uncertainty in our light curves. As previously mentioned, we do not plot $\sigma_{\rm{sys}}$ in Figure 5 because of its largely unconstrained and uninformative behavior. Conversely, Figure 5 shows that $f_{\rm{disk}}$ is the most constrained of these parameters, while the $\alpha_{\rm{dyn}}$ posteriors are generally broader and show minimal preference across individual ejecta fits. The velocity systematics $\beta_{\phi}$ lie somewhere in between, with the Nedora and KruFo GW+EM analyses preferring slower velocities ($\beta_{\phi}>0$) and the KruFo EM and DiCo analyses preferring faster velocities ($\beta_{\phi}<0$). For the velocity values allowed by the range of $\beta_{\phi}$, we find that, compared to the recovered value of $v_{\rm{ej}}=v_{\rm{d}}$ = 0.14 from Figure 3, the KruFo20, DiCo20, and Nedora21 forward models recover relative uncertainties on the velocity of 30.7%, 138%, and 63%. For all but the KruFo20 fit, these relative differences are substantially higher than the 33% originally reported in Ref. [57] from which the ejecta velocity fits originate. Given the narrow peaks of the $f_{\rm{disk}}$ posteriors, we interpret the $f_{\rm{disk}}$ parameter as setting the scale for the wind ejecta mass, with the $\alpha_{\rm{dyn}}$ , $\beta_{\phi}$, and $\sigma_{\rm{sys}}$ parameters allowing our inference enough flexibility to fit the AT2017gfo light curves. Likewise, we expect the ejecta mass scale set by $f_{\rm{disk}}$ to be the dominant parameter determining the light curve behavior (see, e.g., [99]); however, recent studies have shown that velocity profiles can contribute significantly to variations in the light curves [100, 101]. Our velocity systematics can have substantial effects on inference, corroborating the findings of these recent studies. These results motivate more detailed studies of versatile ejecta fits which, among other effects, specifically focus on a separate treatment of dynamical and secular ejecta velocities.
@@
Within the context of no velocity systematics ($\beta_{\phi}=0)$, a closer look at the $R_{\rm{1.4}}$ vs. $\delta$ two-dimensional posterior in Figure 6 indicates a cubic relationship between the two parameters. Specifically, as the mass asymmetry of the binary increases $\delta\rightarrow 0.5$, the radius of the individual stars decreases $R_{\rm{1.4}}\leftarrow 0$, indicating that more asymmetric binaries prefer a softer (more compressible) EOS. This behavior is exactly the same as that observed in Figure 1, where the ejecta velocity contours (dotted lines ranging from 0.1 to 0.2) dictated the $R_{\rm{1.4}}$ vs. $\delta$ relationship. The persistence of this restrictive behavior from the naive contours in Figure 1 to our posteriors in Figure 6 strongly suggests that future analyses of the NS radius and EOS must include both mass and velocity systematic parameters, lest they introduce significant biases on top of the ones already present in selecting an ejecta forward model.
@@
Finally, in the case where the fraction of ejected disk material is fixed to $f_{\rm{disk}}$ $=30\%$, we recover low quality posteriors that clearly struggle to fit the data. Specifically, the wind mass $m_{\rm{w}}$ is unconstrained compared to the dynamical mass $m_{\rm{d}}$, resulting in prohibitively large uncertainties in the light curves.
@@ V Conclusion
In this paper, we have examined multimessenger Bayesian inference of GW170817 and its kilonova AT2017gfo using forward-modeling ejecta predictions which are based on numerical relativity simulations of neutron star mergers. Using a fixed baseline kilonova model [52], we assess the impact of systematic uncertainty in ejecta modeling, using three forward model ejecta fits presented in Refs. [45, 46, 47]. We find that these three forward models are in tension with one another and necessarily over-constrain ejecta inference (i.e., they make more predictions than they have parameters). Using concrete interpretation of GW170817/AT2017gfo, and after adding a generous budget for systematic uncertainty above and beyond stated confidence, we illustrate the impact of the tension in these three models on multimessenger interpretation: we find distinctly different conclusions about the underlying binary with each ejecta model, even allowing for systematics. We furthermore point out that adopting stronger assumptions about ejecta, which omit one or more of our systematic uncertainty parameters, lead to even more extreme conclusions about AT2017gfo/GW170817.
@@
Our study builds upon and extends work by [52], which emphasized purely phenomenological kilonova inference with a fiducial $\sigma_{\rm sys}=0.5$ or a similar prior on $\sigma_{\rm sys}$, without ejecta forward modeling. Except for when we strongly limit our assumptions about ejecta systematics, our inferences about the observed kilonova ejecta are identical.
We highlight that key elements of these models may be used well outside of their region of calibration (e.g., ejecta velocity) when used in conjunction with the surrogate kilonova model from our prior work in Ref. [52]. The primary observations highlighted by our analysis are:
@@
  • Fits to numerical relativity simulations of neutron star mergers which predict mass ejection suffer from systematic uncertainties stemming from the parameter coverage and accuracy of the simulations used to derive them. Unfortunately, multimessenger inference requires confidence in these predictions well outside conventional configurations, to confidently assess extreme scenarios.
@@
  • Given a kilonova model which can provide tight constraints on ejecta parameters, multimessenger inference about binary parameters like the neutron star radius and, consequently, the dense matter equation of state are strongly dependent on the chosen forward model and its ejecta predictions. The ejecta velocity in particular seems to have a strong impact, as ours and other contemporary inferences favor velocities at the low end of calibrated fits.
@@
  • Corroborating previous results (e.g. [100, 101]), we find that a treatment of both mass and velocity systematic uncertainties is required for maximally unbiased inference of ejecta forward model predictions.
@@
  • Within the context of our model, we find that the fraction of the ejected disk lies around $4-16\%$ for a GW170817-like event, with said fraction also dependent on the chosen forward model. This agrees with predictions from GRMHD simulations with full neutrino transport, but is in tension with assumptions typically made in the literature (up to $40\%$ of disk ejection).
@@
Our results suggest that there are currently no broadly applicable prescriptions which accurately map binary parameters to ejecta predictions or correctly assume the ejected fraction of an arbitrary post-merger accretion disk. The lack of such prescriptions strongly motivates future studies which examine the connection between numerical relativity simulations of neutron star mergers and their ejecta predictions. While future multimessenger observations of neutron star mergers will shed more light on the broader population of mergers, in the meantime, our conclusions suggest that significantly more additional numerical relativity studies are necessary to understand the connection between binary and ejecta parameters. Additionally, for kilonova models to be confidently used in multimessenger analyses such as the proof of concept presented in this manuscript, more work is required to improve the fidelity of the radiative transfer simulations used to model these events. Likewise, in understanding these connections, we encourage future studies to carefully consider the contexts within which each of the ejecta forward models is applicable and assume appropriate levels of uncertainty in their analyses.
@@ VI Acknowledgments
ROS and MR acknowledge support from NSF AST 1909534 and 2206321. KW acknowledges support from NSF AST-2319326 and PHY-2012057. ROS also acknowledges support via NSF PHY 2012057, 2309172, and the Simons Foundation. The work by CLF, CJF, MR, MRM, OK, and RTW was supported by the US Department of Energy through the Los Alamos National Laboratory (LANL). This research used resources provided by LANL through the institutional computing program. Los Alamos National Laboratory is operated by Triad National Security, LLC, for the National Nuclear Security Administration of U.S. Department of Energy (Contract No. 89233218CNA000001). This document has been assigned LA-UR-25-22248.

### v2

@@ III.2 Allowing no uncertainty in ejecta velocity ($\alpha_{\rm{dyn}},f_{\rm{disk}}$ free, $\beta_{\phi}=0$)
However, without an additional parameter to encapsulate the ejecta systematics, we find several notable differences, largely dictated by the $v_{\rm{ej}}$ contours in Figure 1. Specifically, we find that for all three ejecta fits, the GW+EM analyses all favor large radii in order to replicate the low dynamical ejecta velocities necessitated by the $v_{\rm{ej}}=v_{d}$ posteriors in Figure 4. Based on Figure 1, a tight constraint on $v_{\rm ej}$ requires either asymmetric NS closer to more conventional radii ($\delta\simeq 0.4$ and $R_{1.4}\simeq 15{\rm km}$) or symmetric NS with extreme radii. Indeed, the union of all three $\delta-R_{1.4}$ posteriors in Figure 6 closely mirrors the underlying $v_{\rm ej}$ contours. As a result, this constraint forces the posterior away from the peak GW likelihood, which favors lower NS radii and more symmetric binaries. In all three cases where we omit systematic error in velocity ($\beta_{\phi}=0$), joint GW+EM inference mildly favors asymmetric binaries $\delta\simeq 0.1-0.2$ with large NS radius $R_{1.4}>15{\rm km}$.
III.3 Allowing no uncertainty in ejecta velocity and disk mass ejection fraction ($\alpha_{\rm{dyn}}$ free, $\beta_{\phi}$ = 0 and $f_{\rm{disk}}=0.30$)
We also consider the case of a fixed disk mass ejection fraction, in which $f_{\rm{disk}}$ is not a free parameter, but restricted to the fixed value $f_{\rm{disk}}$ $30\%$. The purpose of this analysis is to evaluate the behavior of the ejecta fits in predicting light curves under the assumption of a generally applied ejected disk fraction. The top half of Figure 7 displays the same ejecta parameters as in Figure 4. While the dynamical mass $m_{\rm{d}}$ and viewing angle $\theta$ are somewhat consistent with prior results, the wind mass $m_{\rm{w}}$ and dynamical velocity $v_{\rm{d}}$ posteriors yield substantially different results. In the bottom half of Figure 7, we plot the light curves corresponding to the Nedora21 GW+EM posteriors shown in the top part of Figure 7. While the light curves corresponding to the median posterior values (solid lines) are clearly worse fits than those in Figure 5, the more striking feature lies in the exceptionally large uncertainties across all broadband filters. More specifically, the uncertainty bounds are much broader than before, resulting from the broad dynamic range of the $m_{\rm{w}}$ posteriors. The difficulty in constraining $m_{\rm{w}}$ comes from $f_{\rm{disk}}$ being directly tied to the amount of ejected wind mass; as such, what the $m_{w}$ posteriors in Figure 7 are really showing us is the Monte Carlo integrator trying to reconcile the values of $m_{w}$ that actually fit the data with the constricting value of $f_{\rm{disk}}$ $30\%$ that fixes the wind ejecta mass.
@@
All of the samples generated by our analyses, including the GW likelihood data, can be found at [113].
@@ IV Discussion
Unsurprisingly, inference analyses using a joint EM and GW dataset are more constraining than using EM-only data. This behavior is most evident for the $\mathcal{M}_{c}$ and $\delta$ posteriors in Figure 3, which highlight the impact of including the gravitational likelihood $\mathcal{L}_{\rm{GW}}$. The $\mathcal{M}_{c}$ posteriors for each fit clearly converge on a preferred value when supplemented with GW information; likewise, the $\delta$ posteriors become narrower on both global (dynamic range of posterior values) and local (per-ejecta-fit posterior values) scales. With regard to the NS radius $R_{\rm{1.4}}$ posteriors, the difference between EM and GW+EM analyses is less striking. More prevalent, however, are the different regions of the NS radius parameter space covered by the posteriors corresponding to each ejecta fit. Specifically, the GW+EM predictions for $R_{\rm{1.4}}$ are roughly between 8–15 km, 10–20 km, and 10–40 km for the Nedora21  KruFo20  and DiCo20 forward models, respectively. These results strongly imply that, provided a constraining kilonova model, multimessenger inference about neutron star radius and, consequently, the dense matter equation of state is entirely dependent on the chosen forward model and, more importantly, the chosen prescription for mapping BNS merger ejecta to kilonova model input ejecta.
@@
Keeping in mind the individual contributions from the GW and EM information, the joint GW/EM inferences shown in Figure 3 make sense. For almost all parameters, the addition of GW information only slightly perturbs the answers derived from EM information alone. This behavior is most evident in the bottom panel of Figure 3, where narrower priors on $f_{\rm{disk}}$ directly affect the ejected wind mass and thus the EM likelihood, shifting the $\delta$ and $R_{\rm{1.4}}$ posteriors to better agreement across ejecta fits.
@@
As explained in Section II.2, we introduce the $\alpha_{\rm{dyn}}$, $f_{\rm{disk}}$, and $\beta_{\phi}$ free parameters to quantify the systematic uncertainty of the ejecta fit predictions, with $\sigma_{\rm{sys}}$ encompassing all other sources of systematic uncertainty in our light curves. As previously mentioned, we do not plot $\sigma_{\rm{sys}}$ in Figure 3 because of its largely unconstrained and uninformative behavior. Conversely, Figure 3 shows that $f_{\rm{disk}}$ is the most constrained of these parameters, while the $\alpha_{\rm{dyn}}$ posteriors are generally broader and show minimal preference across individual ejecta fits. The velocity systematics $\beta_{\phi}$ lie somewhere in between, with the Nedora and KruFo GW+EM analyses preferring slower velocities ($\beta_{\phi}>0$) and the KruFo EM and DiCo analyses preferring faster velocities ($\beta_{\phi}<0$). For the velocity values allowed by the range of $\beta_{\phi}$, we find that, compared to the recovered value of $v_{\rm{ej}}=v_{\rm{d}}$ = 0.14 from Figure 4, the KruFo20, DiCo20, and Nedora21 forward models recover relative uncertainties on the velocity of 30.7%, 138%, and 63%. For all but the KruFo20 fit, these relative differences are substantially higher than the 33% originally reported in Ref. [60] from which the ejecta velocity fits originate.
Given the narrow peaks of the $f_{\rm{disk}}$ posteriors, we interpret the $f_{\rm{disk}}$ parameter as setting the scale for the wind ejecta mass, with the $\alpha_{\rm{dyn}}$, $\beta_{\phi}$, and $\sigma_{\rm{sys}}$ parameters allowing our inference enough flexibility to fit the AT2017gfo light curves. Likewise, we expect the ejecta mass scale set by $f_{\rm{disk}}$ to be the dominant parameter determining the light curve behavior (see, e.g., [83]); however, recent studies have shown that velocity profiles can contribute significantly to variations in the light curves [114, 115]. Our velocity systematics can have substantial effects on inference, corroborating the findings of these recent studies. These results motivate more detailed studies of versatile ejecta fits which, among other effects, specifically focus on a separate treatment of dynamical and secular ejecta velocities.
@@
Within the context of no velocity systematics ($\beta_{\phi}=0)$, a closer look at the $R_{\rm{1.4}}$ vs. $\delta$ two-dimensional posterior in Figure 6 indicates a cubic relationship between the two parameters. Specifically, as the mass asymmetry of the binary increases $\delta\rightarrow 0.5$, the radius of the individual stars decreases $R_{\rm{1.4}}\rightarrow 0$, indicating that more asymmetric binaries prefer a softer (more compressible) EOS. This behavior is exactly the same as that observed in Figure 1, where the ejecta velocity contours (dotted lines ranging from 0.1 to 0.2) dictated the $R_{\rm{1.4}}$ vs. $\delta$ relationship. The persistence of this restrictive behavior from the naive contours in Figure 1 to our posteriors in Figure 6 strongly suggests that future analyses of the NS radius and EOS must include both mass and velocity systematic parameters, lest they introduce significant biases on top of the ones already present in selecting an ejecta forward model.
@@
Finally, in the case where the fraction of ejected disk material is fixed to $f_{\rm{disk}}$ $30\%$, we recover low quality posteriors that clearly struggle to fit the data. Specifically, the wind mass $m_{\rm{w}}$ is unconstrained compared to the dynamical mass $m_{\rm{d}}$, resulting in prohibitively large uncertainties in the light curves.
@@ V Conclusion
In this paper, we have examined multimessenger Bayesian inference of GW170817 and its kilonova AT2017gfo using forward-modeling ejecta predictions which are based on numerical relativity simulations of neutron star mergers. Using a fixed baseline kilonova model [56] and an established, yet imperfect mapping between BNS merger and kilonova ejecta, we assess the impact of systematic uncertainty in ejecta modeling, using three forward model ejecta fits presented in Refs. [45, 46, 47]. We find that these three forward models are in tension with one another and necessarily over-constrain ejecta inference (i.e., they make more predictions than they have parameters). Using concrete interpretation of GW170817/AT2017gfo, and after adding a generous budget for systematic uncertainty above and beyond stated confidence, we illustrate the impact of the tension in these three models on multimessenger interpretation: we find distinctly different conclusions about the underlying binary with each ejecta model, even allowing for systematics. We furthermore point out that adopting stronger assumptions about ejecta, which omit one or more of our systematic uncertainty parameters, lead to even more extreme conclusions about AT2017gfo/GW170817. The posteriors we infer when omitting ejecta velocity systematics directly challenge the results from previous multimessenger analyses which marginalize over the ejecta velocity entirely [55, 46].
@@
Our study builds upon and extends the work in Ref. [56], which emphasized purely phenomenological kilonova inference with a fiducial $\sigma_{\rm sys}=0.5$ or a similar prior on $\sigma_{\rm sys}$, without ejecta forward modeling. Except for when we strongly limit our assumptions about ejecta systematics, our inferences about the observed kilonova ejecta are identical. We highlight that key elements of these ejecta fits may be used well outside of their region of calibration (e.g., ejecta velocity) when used in conjunction with the surrogate kilonova model from our prior work in Ref. [56].
The primary observations highlighted by our analysis are:
@@
  • Fits to numerical relativity simulations of neutron star mergers which predict mass ejection suffer from systematic uncertainties stemming from the parameter coverage and accuracy of the simulations used to derive them. Unfortunately, multimessenger inference requires confidence in the imperfect mapping between BNS merger and kilonova model ejecta, guiding these predictions well outside conventional configurations to properly assess extreme scenarios. At present, we are unable to confidently and consistently connect binary and ejecta parameters using GW170817/AT2017gfo’s rich observational datasets, thus also dooming future multimessenger inferences to a similar fate.
@@
  • Given a kilonova model which can provide tight constraints on ejecta parameters, multimessenger inference about binary parameters like the neutron star radius and, consequently, the dense matter equation of state are strongly dependent on the chosen forward model and its ejecta predictions which are primarily guided by the chosen prescription for mapping BNS merger ejecta to kilonova model input ejecta. The ejecta velocity in particular seems to have a strong impact, as ours and other contemporary inferences favor velocities at the low end of calibrated fits.
@@
  • Our study examines the inconsistencies between ejecta fits introduced by mapping binary parameters to ejecta mass and velocity, challenging the results from previous multimessenger studies which marginalize over ejecta velocity. Corroborating previous results (e.g. [114, 115]), we find that a treatment of both mass and velocity systematic uncertainties is required for maximally unbiased inference of ejecta forward model predictions.
@@
  • Within the context of our model, we find that the fraction of the ejected disk lies around $4-16\%$ for a GW170817-like event, with said fraction also dependent on the chosen forward model. This agrees with predictions from sub-second GRMHD simulations with full neutrino transport, but is in tension with assumptions typically made in the literature (up to $40\%$ of disk ejection).
  • We emphasize that some of the outcomes and discussed pitfalls in this work are likely to also be related to the specific, established choice of mapping between BNS merger ejecta and kilonova model ejecta inputs. Although conventions exist in the literature, they are incomplete, not one-to-one, and additional work establishing more physical mappings is necessary for accurate multimessenger inference.
@@
Our results suggest that there are currently no broadly applicable prescriptions which accurately map binary parameters to ejecta predictions or correctly assume the ejected fraction of an arbitrary post-merger accretion disk. A large portion of the mapping incompatibility boils down to the difficulty in prescribing the ejecta from numerical relativity simulations to the standard assumption of two distinct, angle-separated kilonova ejecta components. The lack of such prescriptions strongly motivates future studies which examine the connection between numerical relativity simulations of neutron star mergers and their ejecta predictions. While future multimessenger observations of neutron star mergers will shed more light on the broader population of mergers, in the meantime, our conclusions suggest that significantly more additional numerical relativity studies are necessary to understand the connection between binary and ejecta parameters. Additionally, for kilonova models to be confidently used in multimessenger analyses such as the proof of concept presented in this manuscript, more work is required to improve the fidelity of the radiative transfer simulations used to model these events. Likewise, in understanding these connections, we encourage future studies to carefully consider the contexts within which each of the ejecta forward models is applicable and assume appropriate levels of uncertainty in their analyses.
@@ VI Acknowledgments
ROS and MR acknowledge support from NSF AST 1909534 and 2206321. KW acknowledges support from NSF AST-2319326 and PHY-2012057. ROS also acknowledges support via NSF PHY 2012057, 2309172, and the Simons Foundation. The work by CLF, CJF, MR, MRM, OK, and RTW was supported by the US Department of Energy through the Los Alamos National Laboratory (LANL). This research was made possible in part by the Information Science & Technology Institute (ISTI) at LANL. This research used resources provided by LANL through the institutional computing program. Los Alamos National Laboratory is operated by Triad National Security, LLC, for the National Nuclear Security Administration of U.S. Department of Energy (Contract No. 89233218CNA000001). This document has been assigned LA-UR-25-22248.